Direction: Solve each word problem involving a related rates. Make a sketch of your work

if necessary, and then apply differentiation.


A water tank in the form of an inverted cone is being emptied at the rate of 6 cubic

feet per minute. The altitude of the cone is 24 feet and the base radius is 12 feet.

Find how fast the water level is lowering when the water is 10 feet deep.


PLEASE ANSWER MY QUESTION QUICKLY!!

DEADLINE : 05/03/2022 11: 00 PM


1
Expert's answer
2022-05-04T12:38:40-0400

"r\/h=0.5"

"r=0.5h"

"V=1/3 \pi r^2 h=1/3 \pi (0.5h)^2h=1/12 \pi h^3"

"dV/dt=1/4\pi h^2dh/dt"

"-6=1/4 \pi (10)^2 dh/dt"

"dh/dt=\frac{-6}{25\pi}"